OpenStudy (anonymous):

If y varies inversely as x, and y = 18 when x = 2, find y when x = 4. (Points : 4) 9/4 9 36 144

OpenStudy (lgbasallote):

y varies inversely as x so \[y = \frac{k}{x}\] substitute the first condition...y = 18 when x = 2 \[18 = \frac{k}{2}\] \[36 = k\] now solve the second condition...what's y when x = 4 \[y = \frac{k}{x}\] now sub in k and x \[y = \frac{36}{4}\] can you solve it now?

OpenStudy (anonymous):

i think so.

OpenStudy (lgbasallote):

great!! ^_^